# arXiv:2603.00340v1[cs.LG] 27 Feb 2026

### v1

@@ 5.2 Cross-Regional Transferability
Learning from raw trajectories was feasible for all models, including rule-based approaches. However, deep learning architectures such as Transformers achieved state-of-the-art accuracy through inductive transfer and fine-tuning of pretrained models [27]. This transfer-learning approach was particularly advantageous when cross-regional differences in transportation networks and traffic dynamics produced distinct patterns of transportation modes. To evaluate model transferability, we fine-tuned the MOBIS-pretrained models on small subsets of the Geolife dataset to simulate low-shot adaptation. Both SpeedTransformer and LSTM-Attention were fine-tuned on 100 and 200 Geolife trips (approximately 1.1% and 2.2% of the dataset, respectively) for 20 epochs and were subsequently evaluated on the remaining samples. Table 4 reports the overall classification accuracy.
@@ 7 Discussion and Conclusion
Compared to classical and deep ML methods [48; 71; 68; 12; 42], SpeedTransformer outperformed feature-rich models despite using only speed as input. Across different hyperparameter configurations, our model consistently outperformed the LSTM-Attention baseline by 2–3% (Appendix C) and surpassed several other machine learning models by more than 10% (Table 2). It also outperformed the classical rule-based model, which relied solely on speed, by over 30% (Appendix H). This suggests that the Transformer-based model is highly effective at extracting useful information from dense speed sequences that would otherwise underutilized, and it outperforms simpler deterministic model, classical machine-learning models, as well as other deep learning models.
@@
Moreover, this strong performance relied solely on instantaneous speed, without requiring GPS coordinates or additional engineered features. This architectural choice offered several advantages: it mitigates the direct exposure of sensitive location data by decoupling motion dynamics from absolute geographic coordinates[59; 13]. Specifically, we quantify this privacy benefit by contrasting the information entropy required to resolve a spatial position versus that of a speed observation. Because speed is a kinematic scalar inherently decoupled from raw geographic coordinates, it significantly limits the information available to a data-profiling adversary (for detailed discussion see Appendix K. In addition, using only speed as model input simplified preprocessing procedures, thereby enhancing reproducibility across various mobile devices and geographical contexts [7].
@@
In addition, despite the privacy advantages of speed-only inputs, the approach is not entirely immune to sophisticated privacy attacks. Although our model avoids collecting raw coordinates, sequential speed profiles can, in some cases, inadvertently reveal identifiable behavioral patterns [34]. [^4] In our real-world experiment, however, all participants lived in dense urban areas, where individual speed profiles were effectively unrecognizable among hundreds of thousands of co-residents. If an adversary possesses extensive side-channel information, such as high-resolution traffic flow databases or signal timing schedules, they could theoretically attempt to match a specific speed profile to a known corridor, potentially inferring coarse-grained route information. Future work can investigate the integration of enhanced techniques, such as differential privacy and federated learning, to provide formal guarantees against such attacks.

### v2

@@ 5.2 Cross-Regional Transferability
Learning from raw trajectories was feasible for all models, including rule-based approaches. However, deep learning architectures such as Transformers achieved state-of-the-art accuracy through inductive transfer and fine-tuning of pretrained models [Howard and Ruder, 2018]. This transfer-learning approach was particularly advantageous when cross-regional differences in transportation networks and traffic dynamics produced distinct patterns of transportation modes. To evaluate model transferability, we fine-tuned the MOBIS-pretrained models on small subsets of the Geolife dataset to simulate low-shot adaptation. Both SpeedTransformer and LSTM-Attention were fine-tuned on 100 and 200 Geolife trips (approximately 1.1% and 2.2% of the dataset, respectively) for 20 epochs and were subsequently evaluated on the remaining samples. Table 4 reports the overall classification accuracy.
@@ 7 Discussion and Conclusion
Compared to classical and deep ML methods [Ribeiro et al., 2024; Zeng et al., 2023; Yu, 2020; Dabiri et al., 2020; Nawaz et al., 2020], SpeedTransformer outperformed feature-rich models despite using only speed as input. Across different hyperparameter configurations, our model consistently outperformed the LSTM-Attention baseline by 2–3% (Appendix C) and surpassed several other machine learning models by more than 10% (Table 2). It also outperformed the classical rule-based model, which relied solely on speed, by over 30% (Appendix H). This suggests that the Transformer-based model is highly effective at extracting useful information from dense speed sequences that would otherwise underutilized, and it outperforms simpler deterministic model, classical machine-learning models, as well as other deep learning models.
@@
Moreover, this strong performance relied solely on instantaneous speed, without requiring GPS coordinates or additional engineered features. This architectural choice offered several advantages: it mitigates the direct exposure of sensitive location data by decoupling motion dynamics from absolute geographic coordinates[Thompson and Warzel, 2022; De Montjoye et al., 2013]. Specifically, we quantify this privacy benefit by contrasting the information entropy required to resolve a spatial position versus that of a speed observation. Because speed is a kinematic scalar inherently decoupled from raw geographic coordinates, it significantly limits the information available to a data-profiling adversary (for detailed discussion see Appendix K. In addition, using only speed as model input simplified preprocessing procedures, thereby enhancing reproducibility across various mobile devices and geographical contexts [Bolbol et al., 2012b].
@@
In addition, despite the privacy advantages of speed-only inputs, the approach is not entirely immune to sophisticated privacy attacks. Although our model avoids collecting raw coordinates, sequential speed profiles can, in some cases, inadvertently reveal identifiable behavioral patterns [Laureshyn et al., 2009]. [^4] In our real-world experiment, however, all participants lived in dense urban areas, where individual speed profiles were effectively unrecognizable among hundreds of thousands of co-residents. If an adversary possesses extensive side-channel information, such as high-resolution traffic flow databases or signal timing schedules, they could theoretically attempt to match a specific speed profile to a known corridor, potentially inferring coarse-grained route information. Future work can investigate the integration of enhanced techniques, such as differential privacy and federated learning, to provide formal guarantees against such attacks.